OpenStudy (anonymous):

An isosceles triangle has one angle that measures 30 degrees and two angles that each measure 10x+25. Find the value of x that makes the sum of the angles equal 180 degrees. Select one: a. 4 b. 5 c. 6 d. 7

OpenStudy (acxbox22):

30+10x+25+10x+25=180 30+20x+50=180 20x+80=180 20x=100 x=5

OpenStudy (anonymous):

thanks (: